[Hallc_running] [New Logentry] Hall C RC update: 10/25/21 (meeting today at 10am)

mack at jlab.org mack at jlab.org
Mon Oct 25 08:50:02 EDT 2021

Logentry Text:
Quick summary. (For full report, see the RC meeting link below.) 
 We had good production in the last 24 hours, mostly at 40-42muA. Hall C beam current was reduced to 30muA for a couple hours when Hall A was increased from 3 to 6 muA. Beam was off 09:45-13:30 on Sunday for a magnet issue that only impacted Hall C.
 Regarding the Secretary of Energy tour, VIP schedules can change unexpectedly, but the last info we were given was that the accelerator will be off 09:30 Monday morning. Tours wrap up at 13:30.

 Pass change for Hall A is today/Monday, starting after the accelerator is locked up again.

 Day shift only: Try to take 1 hour of Dummy target each day. (a 100% live hour) If we miss it on day shift, I'll give instructions to swing or owl. 
There will be an RC meeting today at 10am: 

bluejeans link: https://bluejeans.com/861439920
RC meeting page: https://hallcweb.jlab.org/wiki/index.php/RC_Meeting,_Monday,_October_25_,_2021


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3932132
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20211025/32053565/attachment.html>

More information about the Hallc_running mailing list